# Let’s Get to the Point: LLM-Supported Planning, Drafting, and Revising of Research-Paper Blog Posts

### v1

@@ 6.3. Study Procedure
The task for the deployment study was to share a blog post with others, which can require significantly higher effort than the controlled lab study, which only required voicing comfort in publicly posting the blog post. To control for the study length and ease of recruitment, we opted for a between-subjects study so that each participant only had to write one blog post. At the same time, we allowed participants who wished to participate in the study twice (N=11/26 participants) to use both tools each time in randomized, counter-balanced order. When analyzing the results we treated them as non-paired data. Eighteen participants completed the baseline condition, and 19 participants completed the treatment condition. For participants who wrote two blog posts, their selected papers were assigned randomly to each condition. Participants had up to six days to complete a research-paper blog post using the assigned tool on their own time. We provided participants with a video tutorial on the assigned tool, a link to the paper about which they would be writing, and a link to the assigned tool. We noted that their aim should be to share the final blog post with others, but if they felt that they could not reach the point of being comfortable with sharing it with others, they could let us know and still be compensated. Participants were required to spend at least 20 minutes using the tool to write the blog post. To better understand how participants utilized the two systems, we recorded their interactions, including prompting, inserting, deleting and replacing text. In addition, we also took periodic snapshots of the drafts whenever they (re)opened the systems in their browsers and every 60 seconds during active usage.
@@
Once participants were satisfied with their draft, participants shared a blog post based on their drafts. We provided instructions on how to create blog posts on the Medium platform,222https://medium.com/ but they were allowed to use other publishing platforms if they preferred. They then emailed us the blog post link and a screenshot or link showcasing evidence that they shared the blog post with others (e.g., tweet, email to co-authors, slack channel message, LinkedIn post). Only 2 of the 26 participants (one under both conditions and one under the treatment condition) alerted us to the fact that they could not reach the point at which they felt comfortable sharing the blog post with others in the allotted time.333We note that some participants indicated that the blog post was part of an experiment when sharing with others, which may have increased their willingness to share it; regardless, they were willing to share a post at the end of the study. Participants then completed the same survey from the lab study regarding their experience with the tool. We note that four participants who we realized had not completed 20 minutes of interaction were asked to do so after completing the survey. The only difference was the removal of a few questions irrelevant to this study: a question regarding how much additional time would be needed to feel comfortable posting the blog post publicly and questions regarding a direct comparison between the two tools. Lastly, for the 15 participants who were willing, the first author engaged them in a 15-minute semi-structured interview to discuss their experience with the tool.
6.4. Results
We did not plan to investigate statistical significance in this between-subjects study, given power analyses indicating that we would need a larger sample size than we could obtain to see a significant result for our metric of Levenshtein distance (power=0.95, alpha=0.05, effect size=0.45; requires 29 participants per condition) or blog post satisfaction (power=0.95, alpha=0.05, effect size=0.98; requires 83 participants per condition). Throughout our analysis of the deployment study logs, we exclude the three study runs in which the participant made no change to the initial draft, as there is no interaction in producing the blog post. We also exclude six study runs in which the participant claimed to have made edits outside the tool beyond small things like grammar, formatting, and adding images; this is to avoid drawing conclusions based on missing data about effort spent. The resulting log data used thus includes 28 study runs (treatment: 14, baseline: 14). The first author conducted inductive thematic analysis (Braun and Clarke, 2006) of the participants’ survey and interview comments.
6.4.1. RQ1: Blog Post Iteration
When given a time constraint in the lab study, we observed that participants made larger changes to the tool’s initial draft with Papers-to-Posts than with the baseline tool. The deployment study allowed us to investigate if this evident decrease in effort could be reflected in a realistic setting and in terms of other metrics of effort (i.e., number of writing actions). We found that participants using Papers-to-Posts took fewer writing actions in order to create one unit of change in Levenshtein distance to the blog post (Figure 9(c)). We also observed, in alignment with the lab study, that Papers-to-Posts led to greater changes in Levenshtein distance for a given amount of active time, which we define as time taking any action in the tool (Figure 9(b)). Therefore, when using Papers-to-Posts as compared to the baseline, participants experienced reduced effort in iterating on the blog post with respect to both writing actions needed and time needed.
Once more, this is despite the fact that participants did not spend substantially more active time using Papers-to-Posts (M=22.00, Q1=17.25,Q3=34.00) than the baseline (M=22.50,Q1=16.00,Q3=27.00) (Appendix 26(a)). Furthermore, we again saw that participants changed the length of their blog posts, measured in characters, less using Papers-to-Posts (M=-5.5, Q1=-525.25, Q3=961.00) than using the baseline (M=710.00, Q1=47.25, Q3=4261.25) (Appendix 26(b)). In fact, Figure 9 shows that the maximum active time and writing actions taken by a baseline participant is higher than that of a treatment participant. This provides some evidence that, for writers who tend to spend more time than average on the task, Papers-to-Posts may help them to reduce the time they spend on it.
Participants’ survey and interview comments reflected this discrepancy between the two tools. With respect to the baseline, participants most commonly cited the initial draft as a helpful aspect of the tool. Nine participants under the baseline condition specifically noted that it was useful as a starting point. Focusing on the initial draft, baseline participants often spoke of editing or adding to it rather than replacing it. For example, P20 explained, ”It’s easier to critique than it is to generate,” and P26 shared, ”It was very useful to have a straw-person of what the blog post could look like (generated by the LLM). I mostly used the draft as a place to edit and steer in the direction I wanted the post to take.”
On the other hand, 15 participants noted the selectable paper bullet points and paragraphs as a helpful feature of Papers-to-Posts. Focusing on the feature designed to generate full blog post sections, participants using Papers-to-Posts often reported doing just that. As an example, P12 commented that the ”outline [to] text generation process is very helpful to get the first draft of the post.” Furthermore, because the tool shows users what bullet points the tool used to generate the initial draft, participants may have been inclined to inspect and iterate upon this selection. This is illustrated by P20, who observed, ”The automatic selection of bullet points to include did not capture most of the information I wanted it to, and so required a lot of manual selection of bullet points, but once I selected those, the generation made it easy to write.” Participants also frequently mentioned the utility of the Papers-to-Posts’s modification abilities. P18 reflected, ”The ”MODIFY WRITING” section is helpful. I like that the interface contains this particular section to support iterating on existing text.” By facilitating modifications to the initial draft, Papers-to-Posts may have encouraged participants to iterate more heavily on the initial draft.
6.4.2. RQ2: Real-World Tool Perceptions
@@
We evaluated participants’ survey results in the same manner as in the lab study (see Section 5.5). With respect to satisfaction with the final blog post quality, we observed no difference in the median satisfaction under Papers-to-Posts(M=6.00,Q1=5.00,Q3=6.00) versus the baseline (M=6.00,Q1=5.25, Q3=6.00) (Figure 5j-m). This discrepancy with the lab study may be related to the fact that participants were not writing under a time constraint in the deployment study. With sufficient time to write the blog post to their liking (only 3 blog posts were not shared with others), participants may not have noticed a pronounced difference in blog post quality between conditions.
Similarly to the lab study, we observed slightly greater satisfaction with the tool under the treatment condition (M=5.60,Q1=5.00, Q3=6.40) than the baseline condition (M=5.40,Q1=4.05,Q3=5.80) (Figure 5n-r). In alignment, four participants who participated twice expressed a preference for Papers-to-Posts, whereas no one shared a preference for the baseline tool. P19 remarked, ”I mean, honestly, I like the first tool better than the second tool… the second tool is a bit simplistic.” However, the complexity and larger associated learning curve of Papers-to-Posts may have decreased its chances of being widely preferred over the baseline. For example, P15 commented, ”It was a lot of information on the interface to deal with. I would prefer less moving things in front of me.” Nevertheless, we once again found slightly lower cognitive load under the treatment condition (M=15.00,Q1=9.50,Q3=17.50) in comparison to the baseline condition (M=15.50,Q1=10.75,Q3=18.75) (Figure 6e-h). This may be related to the fact that, while Papers-to-Posts required understanding a more complex interface, the baseline tool placed greater demands on participants in terms of writing. P24 summarized, ”In this [baseline] version, you have greater autonomy regarding what needs to be done. While this is beneficial, it also requires more effort and planning.”
As shown in Figure 7e-h, participants indicated higher satisfaction with control (treatment: M=5.00,Q1=3.00,Q3=6.00, baseline: M=3.00,Q1=3.00,Q3=4.75) in the treatment condition like in the lab study, but they showed no difference in median satisfaction with respect to perceived control (treatment: M=5.00,Q1=5.00,Q3=6.00, baseline: M=5.00,Q1=3.00,Q3=5.00), perceived ownership (treatment: M=6.00,Q1=4.50,Q3=6.00, baseline: M=6.00,Q1=6.00,Q3=6.00) or satisfaction with ownership (treatment: M=6.00,Q1=4.00,Q3=6.00, baseline: M=6.00,Q1=5.00,Q3=6.00) in the treatment condition. Participants may have been more satisfied with their control of the tool without perceiving a higher control using Papers-to-Posts in the sense that they were comfortable with letting the tool have a certain amount of control in exchange for less effort to generate text.

### v2

@@ 4.2.4. Study Procedure
The task for the deployment study was to share a blog post with others, which can require substantially higher effort than voicing comfort in publicly posting the blog post, as in the lab study. Considering this difference for participant recruitment, we opted for a between-subjects design, so each participant only had to write one blog post. Still, we allowed participants who wished to participate twice (N=11/26 participants) to use both tools in randomized, counter-balanced order. Their selected papers were assigned randomly to each condition. When analyzing these participants’ results, we treated them as unpaired data. Eighteen participants completed the baseline, and 19 completed the treatment. Participants had up to six days to complete a paper blog post using the assigned tool on their own time. We provided a link to the tool, a video tutorial on the tool, and a link to the paper about which to write. We noted that participants should aim to share the final blog post with others, but if they felt that they could not reach that point, they would still be compensated. Participants were required to spend at least 20 minutes using the tool. We recorded participants’ interactions, including prompting and text edits. In addition, we also took snapshots of the blog post whenever they (re)opened the systems in their browsers and every 60 seconds during active usage.
@@
Once participants were satisfied with their draft, they shared it. We provided instructions on how to create blog posts on the Medium platform,121212https://medium.com/ but they were allowed to use other publishing platforms if they preferred. They then emailed us the blog post link and a screenshot or link showcasing evidence that they shared the blog post with others (e.g., tweet, email to co-authors, Slack channel message, LinkedIn post). Only 2 of the 26 participants (one under both conditions and one under the treatment) could not reach the point at which they felt comfortable sharing the blog post with others in the allotted time.131313We note that some participants indicated that the blog post was part of an experiment when sharing with others, which may have increased their willingness to share it; regardless, they were willing to share a post at the end of the study. Participants then completed the same survey from the lab study regarding their experience (see Section 4.1.4).141414We note that four participants who we realized had not completed 20 minutes of interaction were asked to do so after completing the survey. The only difference was the removal of a few questions irrelevant to this study151515a question regarding how much additional time would be needed to feel comfortable posting the blog post publicly and questions regarding a direct comparison between the two tools. Lastly, the first author engaged willing participants (15) in a 15-minute semi-structured interview to discuss their experience and survey responses.
5. Evaluation Studies’ Results
In this section, we present the quantitative and qualitative results of the lab and deployment studies in parallel. To obtain the qualitative results, the first author conducted inductive thematic analysis (Braun and Clarke, 2006) of the participants’ interview and short-answer responses across the two studies. Please note that the hypotheses from the within-subjects lab study were not re-evaluated for significance in the between-subjects deployment study because there were not enough participants to do so (see Section 4.2.1).
5.1. Editing Power
5.1.1. Lab Study - Exploratory
Participants made greater change to the initial LLM draft blog post when using Papers-to-Posts as compared to the baseline tool. For each condition, we calculated the Levenshtein distance between participants’ initial LLM draft blog posts and their blog posts at the point at which they would feel comfortable publicly publishing the blog post. Levenshtein distance measures how many character insertions, deletions, and replacements would be needed to transform one text into another. For statistical testing, we employed a paired-samples t-test and observed that this Levenshtein distance is significantly higher with Papers-to-Posts (M=3977.05, SD=1146.46) compared to the baseline (M=2764.15, SD=1042.16) (Paired-Samples t-Test, t(19)=4.31, p¡.0005).161616We note that this significance test was not part of our initial hypotheses and was therefore exploratory in nature. The difference in Levenshtein distance does not appear attributable to the time spent on the task or the change in blog post length. Participants did not spend substantially more time to reach the point at which they were comfortable publicly publishing the blog post using Papers-to-Posts as compared to the baseline (see Section 5.5). Figure 5a shows Levenshtein distance plotted against the time taken to get to a publishable blog post, with the removal of one outlier. Furthermore, participants changed the length of their blog posts, measured in characters, less with Papers-to-Posts (M=-162.00, Q1=-1126.25, Q3=297.00) than with the baseline (M=1971.00, Q1=965.25, Q3=2514.25) (Appendix 26(b)).171717As you can see in Appendix 26(b), the initial length of treatment drafts was significantly higher than the initial length of baseline drafts.
5.1.2. Deployment Study - RQ1
We investigated if the increase in editing power observed in the lab study would be reflected in a more realistic setting and if Papers-to-Posts required fewer writing actions (i.e., inserting/deleting a character or text span in the draft or LLM instructions)181818We did not count when participants moved, inserted, or deleted a blog post section in Papers-to-Posts as a writing action because these actions involved clicking a button rather than manipulating text. in order to achieve a given amount of change in the blog post. To analyze editing power, we excluded the three study runs in which the participant made no change to the initial draft. We also excluded six study runs in which the participant claimed to have made edits to the blog post outside the tool, beyond small edits like those related to grammar, formatting, and adding images. The resulting analysis thus included 28 study runs (treatment: 14; baseline: 14).
Participants using Papers-to-Posts took fewer writing actions to create one unit of change in Levenshtein distance to the blog post (Figure 5(c)). Papers-to-Posts also led to a greater final Levenshtein distance for a given amount of active time, which we define as time during which participants were taking any actions in the tool (Figure 5(b)). Once more, this is despite the fact that participants did not spend substantially more active time using Papers-to-Posts (M=25.71 minutes, SD=14.72) than the baseline (M=35.93 minutes, SD=34.38) (Appendix 27(a)).191919Considering that there were three outliers with a high amount of active time in the baseline condition, we can also look at the median behavior. We see little difference between the conditions (treatment: M=22.00, Q1=17.25, Q3=34.00; baseline: M=22.50, Q1=16.00,27.00). Furthermore, participants changed the length of their blog posts, measured in characters, less using Papers-to-Posts less using Papers-to-Posts (M=-5.50, Q1=-525.25, Q3=961.00) than using the baseline (M=710.00, Q1=47.25, Q3=4261.25) (Appendix 27(b)).202020As you can see in Appendix 27(b), the initial length of treatment drafts was significantly higher than the initial length of baseline drafts.
5.1.3. Summary
In the lab study, our exploratory analysis revealed that Papers-to-Posts led to increased editing power, in the sense that participants made more change to their blog posts in a given amount of time. In the deployment study, participants experienced the same outcome in a more realistic setting. We further found that participants made more change to their blog posts for a given number of writing actions.
5.2. Blog Post Satisfaction
To measure participants’ satisfaction with their final blog posts, we utilized their Likert-type response to the statement “I am satisfied with the overall quality of the blog post” (Figure 6a,j).
5.2.1. Lab Study - H1
To compare responses under the two conditions in the within-subjects lab study, we employed a non-parametric test, given the data’s Likert nature. Furthermore, we used a sign test rather than a Wilcoxon signed rank test, which is normally used to analyze non-parametric within-subjects data for two conditions of one factor, due to the violation of that test’s assumption of a symmetric distribution of the paired differences about the median. Participants were significantly more satisfied with the final blog post generated with Papers-to-Posts (M=6.00, Q1=5.00, Q3=7.00) as compared to the baseline tool (M=5.00, Q1=5.00, Q3=6.00) (Dependent-Samples Sign Test, S=11, p¡.05). The majority of participants (11) were more satisfied in the treatment condition (equally: 7, less: 2).
Inspecting participants’ Likert responses regarding aspects of their blog post satisfaction, we see that participants particularly demonstrated more satisfaction with the content covered in the Papers-to-Posts blog post (M=6.00, Q1=5.00, Q3=7.00) compared to the baseline blog post (M=5.50, Q1=4.75, Q3=6.00) (Figure 6b). The majority of participants (14) were more satisfied with the content under the treatment condition (equally: 2, less: 4). Although it is not apparent from the medians, participants were overall more satisfied with the blog post structure using Papers-to-Posts (M=6.00, Q1=5.00, Q3=6.25) than the baseline (M=6.00, Q1=5.00, Q3=6.00) (Figure 6d). The plurality of participants (9) were more satisfied with the structure under the treatment (equally: 6, less: 5). Participants did not show increased satisfaction with the blog post’s tone under the treatment (M=6.00, Q1=5.00, Q3=6.00) compared to the baseline (M=5.50, Q1=5.00, Q3=6.00) (Figure 6c). A plurality of participants were not more satisfied with the tone under the treatment (more: 7, equally: 6, less: 7).
Looking at participants’ perceived failure to complete their task (Figure 7c), we observe that participants had lower perceived failure with Papers-to-Posts (M=2.00, Q1=2.00, Q3=3.00) than with the baseline (M=3.00, Q1=2.00, Q3=4.00). The plurality of participants (9) thought that they were more successful in writing a research-paper blog post with Papers-to-Posts compared to the baseline (equally: 7, less: 4). This further corroborates participants’ blog post satisfaction under the treatment condition.
5.2.2. Deployment Study - RQ2a
In the between-subjects deployment study, we observed no difference in the median satisfaction under Papers-to-Posts (M=6.00, Q1=5.00, Q3=6.00) versus the baseline (M=6.00, Q1=5.25, Q3=6.00) in blog post satisfaction (Figure 6j). Again with respect to median satisfaction, participants were equally satisfied with the blog post content (treatment: M=6.00, Q1=5.00, Q3=6.50; baseline: M=6.00, Q1=3.50, Q3=6.00; Figure 6k), tone (treatment: M=6.00, Q1=5.00, Q3=6.00; baseline: M=6.00, Q1=5.00, Q3=7.00; Figure 6l), and structure (treatment: M=6.00, Q1=6.00, Q3=6.00; baseline: M=6.00, Q1=5.25, Q3=7.00; Figure 6m) when using Papers-to-Posts versus the baseline. In terms of perceived failure to complete their task (Figure 7h), participants had equal median perceived failure with Papers-to-Posts (M=3.00, Q1=2.00, Q3=3.00) and the baseline (M=3.00, Q1=2.00, Q3=3.00).
5.2.3. Summary
@@
We observed a significant increase in blog post satisfaction under the treatment condition during the lab study and found that participants were particularly more satisfied with the blog post content, suggesting that Papers-to-Posts best addressed its design goal to support planning blog post content (DG1). We did not observe the same increases in blog post satisfaction in the deployment study. This discrepancy may be related to the fact that participants were not writing under a time constraint in the deployment study. With sufficient time to write the blog post to their liking (only three blog posts were not shared with others), participants may not have noticed a pronounced difference in blog post quality between conditions.
5.3. Tool Satisfaction
For measuring tool satisfaction, we used the five-question Likert scale on tool acceptance from Kocielnik et al. (Kocielnik et al., 2019) (Figure 6e-i). The five responses were averaged to obtain one score.
5.3.1. Lab Study - H2
We used the non-parametric sign test to evaluate the lab study data, as it was Likert data and violated the Wilcoxon signed rank test’s assumption of a symmetric distribution of the paired differences about the median. While participants had a higher median satisfaction with Papers-to-Posts (M=5.90, Q1=5.20, Q3=6.50) than the baseline tool (M=5.60, Q1=4.80, Q3=6.00), this difference was not significant (Dependent-Samples Sign Test, S=13, p=n.s.).
Based on median results, participants under the treatment condition more often found the tool helpful (treatment: M=6.00, Q1=5.75, Q3=7.00; baseline: M=5.50, Q1=4.75, Q3=6.00) and useful for productivity (treatment: M=6.00, Q1=5.00, Q3=7.00; baseline: M=5.50, Q1=5.00, Q3=6.00). Meanwhile, participants were equally likely under both conditions to use the tool if it were available (treatment: M=6.00, Q1=5.00, Q3=7.00; baseline: M=6.00, Q1=5.00, Q3=6.00), recommend the tool (treatment: M=5.50, Q1=4.75, Q3=7.00; baseline: M=5.50, Q1=4.00, Q3=6.00), and find the tool annoying or distracting (treatment: M=2.00, Q1=2.00, Q3=3.00; baseline: M=2.00, Q1=2.00, Q3=3.00). For each of the first four questions, at least four more participants increased rather than decreased their Likert rating from the baseline to the treatment condition. However, only two more participants indicated decreased rather than increased annoyance with Papers-to-Posts (less: 8, more: 6, equal: 6).
5.3.2. Deployment Study - RQ2b
Similarly to the lab study, we observed slightly greater satisfaction with the tool under the treatment condition (M=5.60, Q1=5.00, Q3=6.40) than the baseline condition (M=5.40, Q1=4.05, Q3=5.80) (Figure 6n-r). Based on median results, participants under the treatment condition were more likely to find the tool helpful (treatment: M=6.00, Q1=5.00, Q3=6.50; baseline: M=5.00, Q1=5.00, Q3=6.00), use it if it were available (treatment: M=6.00, Q1=5.00, Q3=7.00; baseline: M=5.00, Q1=4.00, Q3=6.00), and recommend the tool (treatment: M=6.00, Q1=5.00, Q3=6.50; baseline: M=5.00, Q1=3.25, Q3=6.00). They were also less likely to find the tool annoying (treatment: M=2.00, Q1=2.00, Q3=4.00; baseline: M=2.50, Q1=2.00, Q3=4.00). On the other hand, they were equally likely under both conditions to find the tool useful for productivity (treatment: M=6.00, Q1=4.50, Q3=7.00; baseline: M=6.00, Q1=4.00, Q3=6.00).
5.3.3. Summary
Across both studies, we observed higher satisfaction with Papers-to-Posts than the baseline tool, though this difference was not significant according to the lab study results. In particular, participants found Papers-to-Posts more helpful.
5.4. Cognitive Load
To measure cognitive load, we used the NASA TLX Index questionnaire (Hart and Staveland, 1988) and summed the responses for the mental demand, temporal demand, effort, and frustration questions (Figure 7a-b,d-e)212121See 5.2 for discussion of the responses to the performance question.
5.4.1. Lab Study - H3
We utilized the non-parametric sign test to evaluate the lab study data, as it was Likert data and violated the Wilcoxon signed rank test’s assumption of a symmetric distribution of the paired differences about the median. We saw no significant difference between the baseline and treatment conditions (Dependent-Samples Sign Test, S=8, p=n.s.), although participants reported slightly lower cognitive load with Papers-to-Posts (M=13.00, Q1=11.50, Q3=16.30) than the baseline tool (M=14.50, Q1=12.00, Q3=18.30).
We saw the largest difference in median results with respect to perceived effort (treatment: M=3.50, Q1=3.00, Q3=5.00; baseline: M=5.00, Q1=3.75, Q3=5.00). The majority of participants (11) reported using less effort with Papers-to-Posts than with the baseline tool (more: 4, equal: 5). Meanwhile, Papers-to-Posts led to an increase in perceived mental demand (treatment: M=5.00, Q1=3.00, Q3=5.00; baseline: M=4.00, Q1=3.75, Q3=5.25). The small plurality of participants (9) perceived more mental demand with Papers-to-Posts (less: 8, equal: 3). We saw no change in the median result for perceived temporal demand (treatment: M=3.00, Q1=2.00, Q3=4.25; baseline: M=3.00, Q1=2.00, Q3=4.25) or frustration (treatment: M=2.00, Q1=1.00, Q3=4.00; baseline: M=2.00, Q1=2.00, Q3=3.00). An equal number of participants found the treatment more (9) versus less (9) temporally demanding, and the small plurality of participants (9) found the treatment less frustrating (more: 7, equal: 4).
5.4.2. Deployment Study - RQ2c
Once again, we found slightly lower cognitive load under the treatment condition (M=15.00, Q1=9.50, Q3=17.50) in comparison to the baseline condition (M=15.50, Q1=10.75, Q3=18.75) (Figure 7f-g,i-j). Looking at the breakdown of questions, the treatment led to less perceived mental demand (M=4.00, Q1=2.00, Q3=5.50) than the baseline (M=5.00, Q1=3.25, Q3=5.75) but higher perceived effort (treatment: M=5.00, Q1=3.00, Q3=6.00; baseline: M=4.50, Q1=3.00, Q3=5.00). We found no change in the median result for perceived temporal demand (treatment: M=3.00, Q1=2.00, Q3=4.00; baseline: M=3.00, Q1=2.00, Q3=4.00) or frustration (treatment: M=3.00, Q1=2.00, Q3=5.00; baseline: M=3.00, Q1=2.00, Q3=4.75).
5.4.3. Summary
In both studies, we observed a small decrease in cognitive load from the baseline to the treatment condition, but this difference was not significant according to the lab study results. Participants consistently perceived little difference in the tools with respect to temporal demand and frustration. However, their perceptions varied across the studies on whether Papers-to-Posts was more or less mentally demanding and whether it required more or less effort.
5.5. Task Completion Time
5.5.1. Lab Study - H4
We recorded two metrics for task completion time. One was the amount of time participants needed during the session before they would be comfortable publicly publishing the blog post. The other was the perceived amount of time participants would need after the session before they would be comfortable publicly publishing the blog post. As participants reached that point of comfort during most sessions, we focused on the former metric. Because this data was numeric within-subjects data for two conditions of one factor, we used a paired-samples t-test. Participants worked slightly faster with the baseline tool (M=44.32 minutes, SD=12.51) in comparison to Papers-to-Posts (M=45.99 minutes, SD=13.88) (Appendix 26(a)), but this difference was not significant (Paired-Samples t-Test, t(19)=0.55, p=n.s.).
5.6. Fulfillment of Planning Design Goal (DG1)
We analyzed participants’ Likert-type responses to questions related to fulfillment of our first design goal, related to planning the blog post (Figure 8a-b,e-f).
5.6.1. Lab Study - Exploratory
Based on the median results, participants were more confident that all essential information was in the blog post (treatment: M=6.00, Q1=3.00, Q3=6.25; baseline: M=5.00, Q1=3.00, Q3=6.00) and found it easier to incorporate information from the paper (treatment: M=6.00, Q1=5.00, Q3=6.00; baseline: M=5.00, Q1=4.75, Q3=6.00) using Papers-to-Posts (Figure 8a-b). In addition, the interaction logs showed that participants utilized Papers-to-Posts’s DG1-inspired affordance to select relevant paper bullet points and paragraphs to incorporate in the blog post. All participants adjusted the LLM’s pre-selected bullet points to some degree. On average, participants added or removed 28 bullet points and 13 paragraphs. We note that some added bullet points come from automatic selection of bullet points used to generate a new section rather than manual selection. Looking at participants’ logs, when selecting points to generate text, only a few exclusively selected original-text paragraphs rather than bullet points. Participants also provided custom bullet points an average of 1.70 times.
5.6.2. Deployment Study - Exploratory